# Nuclear-physics implications of a subsolar neutron star gravitational-wave observation

### v1

@@ Abstract
If the SSM objects are identified as light NSs, the large tidal effects can be exploited to constrain the NS EoS. This was anticipated in Fig. 1, showing the projected constraints on $\Lambda_{i}$ and $m_{i}$ inferred from an injection using NS2 with a typical O4/O5 SNR. As a proof of principle, we assume an EoS for strange quark matter (SQM3 [38]) compatible with NSs with mass up to $\approx 2M_{\odot}$ (as required by pulsar observations [103]). In the SSM range, this EoS predicts NSs with significantly smaller tidal deformability and radius than ordinary EoS produced by $npe\mu$ matter. Thus, if the event is generated by SSM strange stars, the upper bounds on $\Lambda_{2}$ (i.e. the lighter object) are already tight enough in O4 (and especially O5) to disfavor ordinary EoS at more than $90\%$ confidence level, while the bounds on $\Lambda_{1}$ (i.e. the heavier object) are less constraining.
Notice that the lower bounds Fig. 1 (and Fig. 2, bottom panel) are influenced by a volume effect, since uniform priors on $\Lambda_{i}$ tend to overestimate the posterior at larger values, and the simulated analyses are still prior dependent due to the relatively low SNR. To mitigate this, one could instead adopt log-uniform priors for $\Lambda_{i}$. Our choice remains conservative when estimating the upper bounds.
Distinguishing strange quark stars from ordinary NS supported by soft matter is more challenging, since in this case the predicted smaller $\Lambda_{2}$ can be compatible with the upper bounds. This is the case of the WFF1 EoS [104] and (marginally) of the APR EoS [105], which indeed cannot be excluded solely from the bounds shown in Fig. 1. However, note that the analysis in Fig. 1 is conservative, since the inference is performed considering $\Lambda_{1}$ and $\Lambda_{2}$ as independent parameters while in reality, for a given EoS, they are fixed in terms of the binary masses. To obtain stronger constraints one could perform a Bayesian model selection between two specific EoS models, implementing the EoS-dependent relations $\Lambda_{i}=\Lambda_{i}(m_{i})$ to reduce the number of parameters, as done in Ref. [33] for ordinary NSs. The results of this analysis are presented in Table 3, showing the Bayes factors for two injections (with different mass ratio) assuming APR (or WFF1) EoS and recovery with SQM3, and vice versa. A reference detection in O4 would already decisively exclude APR if the signal was generated by strange stars, whereas the opposite case is less constraining, because the tidal deformability of APR is larger than that of SQM3 and the lower bounds shown in Fig. 1 are not stringent. Remarkably, however, the same putative event detected during O5, or a comparable-mass light SSM binary in O4, would allow distinguishing between APR and SQM3 with exceptionally high statistical confidence. Excluding WFF1 requires higher SNR or light comparable-mass SSM binaries, since in this case the tidal deformability is more similar to that of SQM3. One would need O5 sensitivity (corresponding to SNR = 44) and both SSM components to exclude WFF1 from SQM3 with high confidence.
Overall, we conclude that a putative SSM event detected during LVK O5 run would allow to rule out or confirm the existence of stars made of quarks. This would translate into very stringent constraints on the NS masses and radii, as summarized in Fig. 3, showing a scarcely constrained region of the typical mass-radius diagram of a NS. Since strange stars have significantly smaller radii than ordinary NSs in the SSM range, the constraints on the mass-radius diagram are very informative.
@@
Table 3: Same as Table 2, but for different EoS in NS2. For all the runs, we assume that $m_{2}=0.27M_{\odot}$. The source luminosity distance is $d_{L}=51\,{\rm Mpc}$ ($d_{L}=36\,{\rm Mpc}$) for $m_{1}=0.63M_{\odot}$ ($m_{1}=0.27M_{\odot}$).
Figure 3: The projected mass and radius of subsolar NS obtained from the measurement of the tidal deformability, assuming a subsolar NS binary with SQM3 EoS detected in O4 and O5 with ${\rm SNR}=25$ and ${\rm SNR}=44$, respectively. The $90\%$ contours are obtained by mapping the posteriors of $\Lambda_{i}$ and $m_{i}$ (with injection $m_{1}=0.63M_{\odot}$ and $m_{2}=0.27M_{\odot}$) into those of the radii of the binary components, assuming a relation between $\Lambda$ and the star’s compactness valid for SQM1-3.
@@
Conclusions. We have explored the implications of a putative SSM event detectable by current GW interferometers. Such detection would be a milestone for astrophysics and cosmology, potentially providing the first GW evidence for new physics. We showed that the past O3 LVK sensitivity was insufficient to confidently distinguish between different SSM models. However, the situation will drastically change during the ongoing O4 run and, especially, during the future O5 run. We showed that for a typical SSM event detected in O4, one can confidently identify the nature of the objects (BHs, light NSs, or more exotic BSs). In case data favor a BH interpretation, this would likely imply that PBHs comprise a sizable fraction of dark matter, and that also some stellar-mass mergers in LVK data should be primordial. On the other hand, in case data favor the light NS hypothesis, O5 sensitivity would allow placing stringent upper bounds on the tidal deformability, which could be used to confirm or rule out NSs made of strange quark matter. Overall, our results show that current interferometers have the potential to unveil exciting implications of a SSM detection, and we hope this potential will be fulfilled in the near future.
@@
Acknowledgments. We would like to thank V. De Luca, A. Iovino and A. Riotto for the fruitful discussions. Inference analyses have been performed at the Vera cluster supported by the Italian Ministry of Research and Sapienza University of Rome. This work is partially supported by the MUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity” and by the FARE programme (GW-NEXT, CUP: B84I20000100001). M.V. acknowledges support from the PRIN 2022 grant “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”.
Supplemental Material
A. Parameter estimation
@@
Bayes’ theorem states then that
@@
  • $\mathcal{L}(s|{\bm{\theta}})$ is the probability of having a signal $s$ given the (source) parameters ${\bm{\theta}}$, and is known as the likelihood function; the choice of the likelihood is linked to the noise model that we adopt, for instance, a Gaussian one (see, e.g., [109]);
  • $\pi({\bm{\theta}})$ indicates the prior probability distribution of having the set of parameters ${\bm{\theta}}$; it represents our knowledge about ${\bm{\theta}}$ before we make the measurement; and
@@
We carried out Bayesian inference with the public software BILBY [41], which makes use of DYNESTY nested sampling [110]. To speed up the simulations, we implemented the relative binning technique [111], which greatly speeds up the likelihood evaluation by expanding it over some fiducial parameters. For the real-data part, the relative binning also requires a set of fiducial parameters [112], which are the ones that are meant to maximize the likelihood. Since these are not known a priori, we have obtained maximum likelihood estimates by iterating an optimization algorithm that makes use of routines in the SciPy library. To compute the Bayes factors of different models, we adopted the same broad priors for all inferences (see next section). To speed up the analysis, the only parameter for which we assume a narrow prior is the chirp mass, which is always well-measured.
B. Parameter priors
@@
Table 4: Parameters adopted in the analysis. All priors are defined to be uniform within the reported boundaries, except for $d_{L}$ (uniform in source frame), $\delta$ (cosine), and $\theta_{JN}$ (sine). We further used the trigger time $t_{\rm trigger}=1267725971.0234s$ reported in [43] for SSM200308.
C. Fits for tidal deformabilities
We compile here the fits utilized throughout this work. In deriving the Bayes factors presented in Table 3 and in performing EoS selection with the NS2 model, we connected the tidal deformability to the masses through the approximated relation
@@ C. Fits for tidal deformabilities
To map the posterior into the mass-radius plane, as shown in Fig. 3, we used the following expression to approximate the relationship between a strange star’s compactness and its tidal deformability.
@@
Because the SQM1-3 models only differ by a dimensional scale factor [38], the correspondence between dimensionless quantities such as $C$ and $\Lambda$ remains unaltered when changing from one SQM EoS to another, so that the same coefficients describe the three cases. The accuracy of the fit in the range of masses $\in[0.2,1]M_{\odot}$ is within $0.5\%$.
@@
Table 5: Coefficients in Eq. (10) for the EoS considered in Table 3.

### v2

@@ V.2 NS equation of state
As a proof of principle, as shown in Fig. 1, we assume an EoS for SQM3 compatible with NSs with mass up to $\approx 2M_{\odot}$ (as required by pulsar observations [112]). In the SSM range, this EoS predicts NSs with significantly smaller tidal deformability and radius than ordinary EoS produced by $npe\mu$ matter. Thus, if the event is generated by SSM strange stars, the upper bounds on $\Lambda_{2}$ (i.e. the lighter object) are already tight enough in O4 (and especially O5) to disfavor ordinary EoS at more than $90\%$ confidence level, while the bounds on $\Lambda_{1}$ (i.e. the heavier object) are less constraining. Notice that the lower bounds Fig. 1 (and Fig. 2, bottom panel) are influenced by a volume effect, since uniform priors on $\Lambda_{i}$ tend to overestimate the posterior at larger values, and the simulated analyses are still prior dependent due to the relatively low SNR. To mitigate this, one could instead adopt log-uniform priors for $\Lambda_{i}$. Our choice remains conservative when estimating the upper bounds.
@@
Table 3: Same as Table 2, but for different EoS in NS2. The source luminosity distance is $d_{L}=51\,{\rm Mpc}$ ($d_{L}=36\,{\rm Mpc}$) for $m_{1}=0.63M_{\odot}$ ($m_{1}=0.27M_{\odot}$). For all the runs, we assume that $m_{2}=0.27M_{\odot}$.
Distinguishing strange quark stars from ordinary NS supported by soft matter is more challenging, since in this case the predicted smaller $\Lambda_{2}$ can be compatible with the upper bounds. This is the case of WFF1 and marginally of APR, which indeed cannot be excluded solely from the bounds shown in Fig. 1. However, note that the analysis in Fig. 1 is conservative since the inference is performed considering $\Lambda_{1}$ and $\Lambda_{2}$ as independent parameters, while in reality, for a given EoS, they are fixed in terms of the binary masses. To obtain stronger constraints, one could perform a Bayesian model selection between two specific EoS models, implementing the EoS-dependent relations $\Lambda_{i}=\Lambda_{i}(m_{i})$ in the waveform to reduce the number of parameters, as done in Ref. [36] for ordinary NSs. The results of this analysis are presented in Table 3, showing the Bayes factors for two injections (with different mass ratios) assuming APR or WFF1 EoS and recovery with SQM3 and vice versa. Also, in this case, the results were obtained assuming NS2 but would be very similar for the less conservative model NS1. A reference detection in O4 would already decisively exclude APR if the signal was generated by strange stars. However, the opposite case is less constraining, because the tidal deformability of APR is larger than that of SQM3, and the lower bounds shown in Fig. 1 are not stringent. Remarkably, however, the same putative event detected during O5, or a comparable-mass light SSM binary in O4, would allow distinguishing between APR and SQM3 with exceptionally high statistical confidence. Excluding WFF1 requires higher SNR or light comparable-mass SSM binaries, since in this case the tidal deformability is more similar to that of SQM3. One would need O5 sensitivity (corresponding to SNR = 44) and sufficiently light SSM components to exclude WFF1 from SQM3 with high confidence.
Overall, we conclude that a putative SSM event detected during the LVK O5 run would allow ruling out or confirming the existence of stars made of quarks, a key result toward the quest for the EoS of high-density nuclear matter [41, 42, 43]. This would translate into very stringent constraints on the NS masses and radii, as summarized in Fig. 3, showing a scarcely constrained region of the typical mass-radius diagram of a NS (since there are currently no observational constraints on the radius of subsolar NSs). Since strange stars have significantly smaller radii than ordinary NSs in the SSM range, the constraints on the mass-radius diagram are very informative to discriminate among different EoS.
VI Conclusions
@@
We have explored the implications of a putative SSM event detectable by current GW interferometers. Such detection would be a milestone for astrophysics and cosmology, potentially providing the first GW evidence for new physics. We showed that the past O3 LVK sensitivity was insufficient to confidently distinguish between different SSM models. However, the situation will drastically change during the ongoing O4 run and, especially, during the future O5 run. We showed that for a typical SSM event detected in O4, one can confidently identify the nature of the objects, being BHs, light NSs, or more exotic BSs. In case data favor a BH interpretation, this would likely imply that PBHs comprise a sizable fraction of dark matter and that also some stellar-mass mergers in LVK data should be primordial. On the other hand, in case data favor the light NS hypothesis, O5 sensitivity would allow placing stringent upper bounds on the tidal deformability, which could be used to confirm or rule out NSs made of strange quark matter. Overall, our results show that current interferometers have the potential to unveil exciting implications of a SSM detection, and we hope this potential will be fulfilled in the near future.
Figure 3: The projected mass and radius of a subsolar mass NS obtained from the measurement of the tidal deformability. We use $m_{1}=0.63M_{\odot}$ and $m_{2}=0.27M_{\odot}$ with the SQM3 EoS detected in O4 and O5 with SNR = 25 and SNR = 44, respectively. Assuming a relationship for $\Lambda$ and the star’s compactness valid for SQM1-3, we obtain the 90% contours by mapping the posteriors of $\Lambda_{i}$ and $m_{i}$ into those of the radii of the binary components.
Acknowledgments
@@
We would like to thank V. De Luca, A. Iovino and A. Riotto for the fruitful discussions. Inference analyses have been performed at the Vera cluster supported by the Italian Ministry of Research and Sapienza University of Rome. This work is partially supported by the MUR PRIN Grant No. 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity” and by the FARE programme (GW-NEXT, CUP: B84I20000100001). M.V. acknowledges support from the PRIN 2022 Grant No. “GUVIRP - Gravity tests in the UltraViolet and InfraRed with Pulsar timing”.
Inference simulations have been carried out with BILBY [113], a python-based software for gravitational wave inference. The manuscript content has been derived using publicly available software: matplotlib, corner, json, numpy [114, 115, 116, 117]. Codes are available upon request and can be requested from francesco.crescimbeni@uniroma1.it.
Appendix A Parameter estimation
@@
Bayes’s theorem states then that
@@
  • $\mathcal{L}(s|{\bm{\theta}})$ is the probability of having a signal $s$ given the (source) parameters ${\bm{\theta}}$ and is known as the likelihood function; the choice of the likelihood is linked to the noise model that we adopt, for instance, a Gaussian one (see, e.g., Ref. [121]).
  • $\pi({\bm{\theta}})$ indicates the prior probability distribution of having the set of parameters ${\bm{\theta}}$; it represents our knowledge about ${\bm{\theta}}$ before we make the measurement.
@@
We carried out Bayesian inference with the public software BILBY [44], which makes use of DYNESTY nested sampling [122]. To speed up the simulations, we implemented the relative binning technique [123], which greatly speeds up the likelihood evaluation by expanding it over some fiducial parameters. For the real-data part, the relative binning also requires a set of fiducial parameters [124], which are the ones that are meant to maximize the likelihood. Since these are not known a priori, we have obtained maximum likelihood estimates by iterating an optimization algorithm that makes use of routines in the SciPy library. To compute the Bayes factors of different models, we adopted the same broad priors for all inferences (see the next section). To speed up the analysis, the only parameter for which we assume a narrow prior is the chirp mass, which is always well-measured.
Appendix B Parameter priors
@@
Table 4: Parameters adopted in the analysis. All priors are defined to be uniform within the reported boundaries, except for $d_{L}$ (uniform in source frame), $\delta$ (cosine), and $\theta_{JN}$ (sine). We further used the trigger time $t_{\rm trigger}=1267725971.0234s$ reported in Ref. [47] for SSM200308.
Appendix C Fits for tidal deformabilities
We compile here the fits utilized throughout this work. In deriving the Bayes factors presented in Table III and in performing EoS selection with the NS2 model, we connected the tidal deformability to the masses through the approximated relation
@@ Appendix C Fits for tidal deformabilities
To map the posterior into the mass-radius plane, as shown in Fig. 3, we used the following expression to approximate the relationship between a strange star’s compactness and its tidal deformability:
@@
Because the SQM1-3 models only differ by a dimensional scale factor [41], the correspondence between dimensionless quantities such as $C$ and $\Lambda$ remains unaltered when changing from one SQM EoS to another so that the same coefficients describe the three cases. The accuracy of the fit in the range of masses $\in[0.2,1]M_{\odot}$ is within $0.5\%$.
@@
Table 5: Coefficients in Eq. (10) for the EoS considered in Table III.